OpenStudy (anonymous):

write an equation for x is less than of equal to -2.

sam (.sam.):

\[x \le-2\]

OpenStudy (anonymous):

\[x \ge-2\]\[x \le-2\]

OpenStudy (anonymous):

i got 

OpenStudy (anonymous):

the arrowhead points to the small numbers (< less than) and viceversa